# The expected kinematic matter dipole is robust against source evolution

### v1

@@ 1 Introduction
Long before it was robustly possible in practice, Ellis & Baldwin (1984) (hereafter EB84) proposed a simple test of the CP via the dipole anisotropy in the projected number density of flux-limited catalogues of sources with power-law spectra, $S\propto\nu^{-\alpha}$, and integral source counts per unit solid angle, $\mathrm{d}N/\mathrm{d}\Omega(S>S_{*})\propto S_{*}^{-x}$. A moving observer would see the source flux density to be Doppler boosted up or down, above or below the flux limit $S_{*}$, such that sources would move into or out of the sample, depending on their location on the sky with respect to the boost direction. This significantly enhances the usual aberration ($\propto\beta$) which displaces the sources towards the direction of motion. Consequently the observer moving with respect to the CMB frame would see a matter dipole aligned with the CMB dipole but with amplitude
@@
where $f(r)$ is the normalised distribution of sources in the catalogue to be defined below. DB22 proved that this expression is equivalent to the formulations that require knowledge of the source luminosity function. By observing that Eq. (2) does not obviously reduce to Eq. (1), they cast doubt on the significance of the result obtained by Secrest et al. (2021, 2022), and called for the determination of the luminosity function and its evolution as the only viable approach to correctly predict the expected kinematic matter dipole amplitude.
@@
While clear, its rationale might not be obvious. Here, we explicitly demonstrate that Ellis & Baldwin’s insight was indeed correct. We clarify the issue by first describing exactly how $\tilde{x}$ and $\tilde{\alpha}$ are determined in practice, and then for the first time show that, even given arbitrary source evolution, Eqs. (1) and (2) are perfectly equivalent.
@@ 2.1 The magnification bias $x$
We follow the definitions in DB22. In each shell of $r$ we find a distribution of sources with observed spectral flux density $S>S_{*}$,
$\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})=kS_{*}^{-x(r)},$ (4)
with normalisation $k$, where $x(r)>0$. We then define the normalised distribution
@@
(We assume that the integrals converge, either by appropriate choice of $x(r)$ and/or by additionally introducing an appropriate window function $w(r)$.) This distribution is not usually known for a given source sample and requires the knowledge of its luminosity function. In contrast, the flux distribution integrated over comoving distance is an observed quantity
$\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})=\int_{0}^{\infty}{\rm d}r\,\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})=k\int_{0}^{\infty}{\rm d}r\,S_{*}^{-x(r)}.$ (6)
Note that the observed integral source counts need not exhibit perfect power-law behaviour at all thresholds $S_{*}$. In fact, if $x(r)$ is not a constant then a change in slope is inevitable. Consider e.g. $x(r)$ linear in $r$, i.e. $x(r)=x_{0}+x_{1}r$. Eq. (6) then gives
$\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})=\frac{k}{x_{1}\ln S_{*}}S_{*}^{-x_{0}},\quad{\rm for}\,x_{1}\ln S_{*}>0,$ (7)
where the departure from a pure power-law with index $x_{0}$ is driven by the additional factor $1/\ln S_{*}$. The EB84 test requires the determination of the slope only at the flux density threshold $S_{*}$, as sources with flux densities far from it will not contribute to the dipole amplitude, e.g. a source with very high $S\gg S_{*}$ will remain in the catalogue regardless of the boost applied.
@@
$\begin{split}\tilde{x}\equiv&-\frac{{\rm d}}{{\rm d}\ln S_{*}}\left[\ln\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})\right]\\ =&-\frac{S_{*}}{\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})}\frac{{\rm d}}{{\rm d}S_{*}}\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})\\ =&-\frac{S_{*}}{k\int_{0}^{\infty}{\rm d}r\,S_{*}^{-x(r)}}\frac{{\rm d}}{{\rm d}S_{*}}\left[k\int_{0}^{\infty}{\rm d}r\,S_{*}^{-x(r)}\right]\\ =&-\frac{S_{*}}{k\int_{0}^{\infty}{\rm d}r\,S_{*}^{-x(r)}}\left[-k\int_{0}^{\infty}{\rm d}r\,x(r)S_{*}^{-x(r)-1}\right]\\ =&\frac{1}{\int_{0}^{\infty}{\rm d}r\,\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})}\int_{0}^{\infty}{\rm d}r\,\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})x(r)\\ =&\int{\rm d}r\,f(r)x(r)=x_{\rm eff}.\end{split}$ (8)
This shows that $x$ measured locally at $S_{*}$ in the observed integral number counts equals the appropriately weighted average as given by $x_{\rm eff}$. E.g. for $x(r)$ linear in $r$ we find $\tilde{x}=x_{\rm eff}=x_{0}+{1/\ln S_{*}}$.
@@ 2.2 The spectral index $\alpha$
Rather than integral source counts we require knowledge of differential source counts that can be evaluated at a given flux density $S$. The integral and differential source counts are related as
@@
Using Eq. (4) we find
$\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r{\rm d}S}(r,S)=kx(r)S^{-x(r)-1},$ (10)
where we made use, as did DB22, of the fact that $x(r)$ does not depend on $S$. This can be used to define a second distribution function over comoving distance,
$\begin{split}\tilde{f}(r)\equiv&\left[\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r{\rm d}S}(r,S)\middle/\int_{0}^{\infty}{\rm d}r\,\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r{\rm d}S}(r,S)\right]_{S=S_{*}}\\ =&\,x(r)\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})\bigg{/}\int_{0}^{\infty}{\rm d}r\,x(r)\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*}),\end{split}$ (11)
where we have used Eqs. (4) and (10). Now $\tilde{\alpha}$ is computed as the average spectral index of sources at the threshold $S_{*}$:
@@
which is generally different from $\alpha_{\rm eff}$ (3). E.g. for linear $x(r)$ and linear $\alpha(r)=\alpha_{0}+\alpha_{1}r$, we find
$\tilde{\alpha}=\frac{1}{\tilde{x}}\left(x_{0}\alpha_{0}+\frac{x_{0}\alpha_{1}}{x_{1}\ln S_{*}}+\frac{\alpha_{0}}{\ln S_{*}}+\frac{2\alpha_{1}}{x_{1}(\ln S_{*})^{2}}\right)$ (13)
to be compared with
$\alpha_{\rm eff}=\alpha_{0}+\frac{1}{x_{1}\ln S_{*}}\alpha_{1}.$ (14)
@@ 2.3 Dipole amplitude
using which we now compare Eq. (1) with Eq. (2). Due to the normalisation of $f(r)$ the first terms are trivially equal. The second terms are equal according to Eq. (8). The crucial term therefore is the last:
@@
Therefore the dipole amplitude is determined entirely by the observed quantities, $\tilde{x}$ and $\tilde{\alpha}$, which account implicitly for all source evolution effects. Moreover, as shown by DB22, there is exact equivalence between the dipole amplitude computed using the luminosity function (Maartens et al., 2018; Nadolny et al., 2021; Dalang & Bonvin, 2022; Guandalin et al., 2023), and that obtained by taking $x$ and $\alpha$ to be $r$-dependent (Dalang & Bonvin, 2022). Thus we have also demonstrated that the theoretically expected kinematic matter dipole amplitude does not require any knowledge of the luminosity function.
@@ 3 Discussion
We have shown that the kinematic matter dipole can be robustly predicted using only observed quantities, as per the EB84 prescription (1), which turned out to be fully equivalent to adding up the contributions (2) from shells of comoving distance. Whereas DB22 highlighted the explicit connection between variations in the intrinsic sample properties $x$ and $\alpha$, and the corresponding computation in a fully general relativistic setting (e.g. Challinor & Lewis, 2011), according to their own calculation (see eq. 36), the equivalence we have shown also implies equivalence with the kinematic matter dipole amplitude calculated using the luminosity function. This makes the determination of the luminosity function redundant in this context, vindicating the assertion of EB84 that “source evolution” is irrelevant in the computation of the kinematic matter dipole.
@@
The difficulty in determining the luminosity function was nevertheless emphasised as a major “theoretical systematic” in predicting the kinematic matter dipole (Guandalin et al. (2023), see also Maartens et al. (2018). This was emphasised by DB22 (see also Dalang & Bonvin, 2023) who questioned the significance of the anomaly found by Secrest et al. (2021) by showing that a range of dipole amplitudes is predicted depending on the choice of the functional form of the magnification bias $x(z)$ and evolution bias $b_{\mathrm{e}}(z)$. Of course this uncertainty only arises due to not knowing the exact functional forms that derive directly from the (unknown) luminosity function of the source sample. However as we have shown, none of this is relevant here. Such uncertainties are entirely bypassed in ensuring to use solely the observed quantities $\tilde{x}$ and $\tilde{\alpha}$ in the EB84 test, as is both appropriate and adequate in the present context. Nevertheless the relation we have found between $\tilde{x}$ and $x(r)$ may prove useful in constraining the magnification bias via measurements of the integral source counts, without reference to the (uncertain) luminosity function as has been done by e.g. Wang et al. (2020).
@@
The correlations first suspected between $x(r)$ and $\alpha(r)$ by DB22 are based on a definition of $\alpha$, namely $\alpha_{\rm eff}$, that does not feature in the EB84 test. These correlations have been investigated empirically by Dam et al. (2023) who invoked the principle of maximal entropy to bound their maximum effect to be below 17% in the quasar sample studied by Secrest et al. (2021). However, we have seen that such correlations need not be studied at all in order to obtain a robust estimate of the dipole amplitude defined in terms of $\tilde{\alpha}$.
Thus no alterations to the EB84 dipole prediction are needed to take into account source evolution effects, as long as the power-law assumptions for the integral source counts and for the source spectra hold at the flux limit $S_{*}$. (This had not been necessary to state explicitly by EB84 who considered perfect power-laws throughout.) Possible curvature in integral source counts ${\rm d}N/{\rm d}\Omega(S>S_{*})$ may warrant small corrections to $\tilde{x}$, as illustrated by the difference between the blue and the black curves in Fig. 1. Also the domain over which to integrate $\alpha$, illustrated by the white shaded area in Fig. 2, offers scope for such corrections. While both are expected to be small corrections to an already small effect of $\mathcal{O}(\beta)$, they can be estimated directly from data. Related corrections were proposed by Tiwari et al. (2014) and implemented by Siewert et al. (2021). While they did not recognise the adequacy of performing the fit of $\tilde{x}$ only locally—e.g.  Siewert et al. (2021) say their fit “extends for each survey over one decade”—they demonstrated the feasibility of the procedure, albeit with little effect on their conclusions.
@@
Lastly, the possible correlation between $\alpha$ and $S$ at the flux limit $S_{*}$ does play a role in the EB84 test, as can be seen by studying $\tilde{\alpha}$ in Fig. 2. This provides justification for the procedure employed by Secrest et al. (2021) to simulate mock skies in order to evaluate the significance with which the dipole measurement rejects the null hypothesis (1). First, $N$ points were distributed randomly on the sky in line with the assumption of isotropy. Each point then was randomly assigned a pair of values $S$ and $\alpha$, sampled from the empirical distribution functions ${\rm d}N/{\rm d}S$ and ${\rm d}N/{\rm d}\alpha$. This respects possible correlations between $S$ and $\alpha$ in the data; in the above context, it amounts to ensuring that in each of the simulations, $\tilde{\alpha}$ is conserved. Later, Secrest et al. (2022) found that the results are nearly unchanged even when $S$ and $\alpha$ are picked independently, which they interpreted as implying that there is no significant source evolution. In light of the present work, we can now confirm that this intuition was correct.

### v2

@@ 1 Introduction
Long before it was robustly possible in practice, Ellis & Baldwin (1984) (hereafter EB84) proposed a simple test of the CP via the dipole anisotropy in the projected number density of flux-limited catalogues of sources with power-law spectra, $S\propto\nu^{-\alpha}$, and integral source counts per unit solid angle, $\mathrm{d}N/\mathrm{d}\Omega(S>S_{*})\propto S_{*}^{-x}$. A moving observer would see the source flux density to be Doppler boosted up or down, above or below the flux limit $S_{*}$, such that sources would move into or out of the sample, depending on their flux and their location on the sky with respect to the boost direction, while the usual aberration ($\propto\beta$) simultaneously displaces the sources towards the direction of motion. Consequently the observer moving with respect to the CMB frame would see a matter dipole aligned with the CMB dipole but with amplitude
@@
where $f(r)$ is the normalised distribution of sources in the catalogue to be defined below. DB22 proved that this expression is equivalent to the formulations that require knowledge of the sources’ luminosity function. By observing that Eq. (2) does not obviously reduce to Eq. (1), they cast doubt on the significance of the result obtained by Secrest et al. (2021, 2022), and called for the determination of the luminosity function and its evolution as the only viable approach to correctly predict the expected kinematic matter dipole amplitude.
@@
While clear, its rationale might not be obvious. Here, we explicitly demonstrate that Ellis & Baldwin’s insight was indeed correct. We clarify the issue by first describing exactly how $\tilde{x}$ and $\tilde{\alpha}$ are determined in practice, and then show that, even given arbitrary source evolution, Eqs. (1) and (2) are perfectly equivalent.
@@ 2.1 The magnification bias $x$
We follow the notation in DB22, but here, for completeness, state the general and implicit dependences $x(r)=x(r,S_{*})$ and $f(r)=f(r,S_{*})$. In each shell of $r$ we find a distribution of sources with observed spectral flux density $S>S_{*}$,
$\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*}),$
that decreases with $S_{*}$ such that a power-law index $x$ is defined locally as
$\begin{split}x(r)\equiv&-\frac{\partial}{\partial\ln S_{*}}\left[\ln\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})\right]\\ =&-\frac{S_{*}}{\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})}\frac{\partial}{\partial S_{*}}\left[\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})\right].\end{split}$ (4)
This is the magnification bias. We also define the normalised distribution
@@
This distribution is not usually known for a given source sample and requires the knowledge of its luminosity function, rather than being directly observable. In contrast, the flux distribution integrated over comoving distance, aka the integral source counts, is an observed quantity
$\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})=\int_{0}^{\infty}{\rm d}r\,\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})$ (6)
The EB84 test requires the determination of the integral source counts’ slope only at the flux density threshold $S_{*}$, as sources with flux densities far from it will not contribute to the dipole amplitude, e.g. a source with very high $S\gg S_{*}$ will remain in the catalogue regardless of the boost applied.
@@
$\begin{split}\tilde{x}\equiv&-\frac{\partial}{\partial\ln S_{*}}\left[\ln\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})\right]\\ =&-\frac{S_{*}}{\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})}\frac{\partial}{\partial S_{*}}\left[\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})\right]\\ =&-\frac{S_{*}}{\frac{{\rm d}N}{{\rm d}\Omega}(S>S_{*})}\int_{0}^{\infty}{\rm d}r\,\frac{\partial}{\partial S_{*}}\left[\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})\right]\\ =&\frac{1}{\int_{0}^{\infty}{\rm d}r\,\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})}\int_{0}^{\infty}{\rm d}r\,\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})x(r)\\ =&\int{\rm d}r\,f(r)x(r)=x_{\rm eff},\end{split}$ (7)
where in the third line we used Eq. 4. This shows that $x$ measured locally at $S_{*}$ in the observed integral number counts equals the appropriately weighted average as given by $x_{\rm eff}$.
@@ 2.2 The spectral index $\alpha$
Rather than integral source counts we require knowledge of differential source counts that can be evaluated at a given flux density $S$. The integral and differential source counts are related by
@@
This can be used to define a second normalised distribution function over comoving distance, $\tilde{f}(r)=\tilde{f}(r,S_{*})$,
$\tilde{f}(r)\equiv$ $\left[\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r{\rm d}S}(r,S)\middle/\int_{0}^{\infty}{\rm d}r\,\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r{\rm d}S}(r,S)\right]_{S=S_{*}}$
$=$ $\,\frac{{\rm d}}{{\rm d}S_{*}}\left[\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})\right]\bigg{/}\int_{0}^{\infty}{\rm d}r\,\frac{{\rm d}}{{\rm d}S_{*}}\left[\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})\right]$
$=$ $\,x(r)\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*})\bigg{/}\int_{0}^{\infty}{\rm d}r\,x(r)\frac{{\rm d}N}{{\rm d}\Omega{\rm d}r}(r,S>S_{*}),$ (9)
where we have used Eqs. (8) and again (4). Now $\tilde{\alpha}$ is computed as the average spectral index of sources at the threshold $S_{*}$,
@@
which is generally different from $\alpha_{\rm eff}$ (3).
@@ 2.3 Dipole amplitude
using which we now compare Eq. (1) with Eq. (2). Due to the normalisation of $f(r)$ their first terms are trivially equal. Their second terms are equal according to Eq. (7). The crucial term therefore is the last:
@@
Therefore the dipole amplitude is determined entirely by the observed quantities, $\tilde{x}$ and $\tilde{\alpha}$, which account implicitly for all source evolution effects. Moreover, as shown by DB22, in the absence of boundary terms as is appropriate for samples with smoothly decaying redshift distributions (see Discussion), there is exact equivalence between the dipole amplitude computed using the luminosity function (Maartens et al., 2018; Nadolny et al., 2021; Dalang & Bonvin, 2022; Guandalin et al., 2023), and that obtained by taking $x$ and $\alpha$ to be $r$-dependent (Dalang & Bonvin, 2022). Thus we have also demonstrated that the theoretically expected kinematic matter dipole amplitude does not require any knowledge of the luminosity function.
@@ 3 Discussion
We have shown that the kinematic matter dipole can be robustly predicted using only observed quantities, as per the EB84 prescription (1), which turned out to be fully equivalent to adding up the contributions (2) from shells of comoving distance. Whereas DB22 highlighted the explicit connection between variations in the intrinsic sample properties $x$ and $\alpha$, and the corresponding computation in a fully general relativistic setting (e.g. Challinor & Lewis, 2011), according to their own calculation (see eq. 36), the equivalence we have shown also implies equivalence with the kinematic matter dipole amplitude calculated using the luminosity function. This makes the determination of the luminosity function redundant in this context, vindicating the assertion of EB84 that source evolution is irrelevant in the computation of the kinematic matter dipole.
@@
The difficulty in determining the luminosity function was nevertheless emphasised as a major “theoretical systematic” in predicting the kinematic matter dipole (Guandalin et al. (2023), see also Maartens et al. (2018)). This was first illustrated by DB22 (see also Dalang & Bonvin, 2023) who questioned the significance of the anomaly found by Secrest et al. (2021) by showing that a range of dipole amplitudes is predicted depending on the choice of the functional form of the magnification bias $x(z)$ and evolution bias $b_{\mathrm{e}}(z)$. This uncertainty only arises due to not knowing the exact functional forms that derive directly from the (unknown) luminosity function of the source sample. However as we have shown, none of this is relevant here. Such uncertainties are entirely bypassed in ensuring to use solely the observed quantities $\tilde{x}$ and $\tilde{\alpha}$ in the EB84 test, as is both appropriate and adequate in the present context. Nevertheless the relation we have found between $\tilde{x}$ and $x(r)$ may prove useful in constraining the magnification bias via measurements of the integral source counts, without reference to the (uncertain) luminosity function as has otherwise been demonstrated by e.g. Wang et al. (2020).
@@
The correlations first suspected between $x(r)$ and $\alpha(r)$ by DB22 are based on a definition of $\alpha$, namely $\alpha_{\rm eff}$, that does not feature in the EB84 test. These correlations have been investigated empirically by Dam et al. (2023) who invoked the principle of maximal entropy to bound their maximum effect to be below 17% in the quasar sample studied by Secrest et al. (2021). As we have seen that such correlations need not be studied at all in order to obtain a robust estimate of the dipole amplitude defined in terms of $\tilde{\alpha}$.
@@
Instead, the possible correlation between $\alpha$ and $S$ at the flux limit $S_{*}$ does play a role in the EB84 test, as can be seen by studying $\tilde{\alpha}$ in Fig. 2. This provides justification for the procedure employed by Secrest et al. (2021) to simulate mock skies in order to evaluate the significance with which the dipole measurement rejects the null hypothesis (1). First, $N$ points were distributed randomly on the sky in line with the assumption of isotropy. Each point then was randomly assigned a pair of values $S$ and $\alpha$, sampled from the empirical distribution function ${\rm d}N/{\rm d}S{\rm d}\alpha$. This respects possible correlations between $S$ and $\alpha$ in the data; in the above context, it amounts to ensuring that in each of the simulations, $\tilde{\alpha}$ is conserved. Later, Secrest et al. (2022) found that the results are nearly unchanged even when $S$ and $\alpha$ are picked independently, which they interpreted as implying that there is no significant source evolution. In light of the present work, we can now confirm that this intuition was correct.
In summary, for samples whose redshift distributions fall sufficiently smoothly towards both low and high redshifts, no alterations to the EB84 dipole prediction are needed to take into account source evolution effects, as long as the power-law assumptions for the integral source counts and for the source spectra hold at the flux limit $S_{*}$. (This had not been necessary to state explicitly by EB84 who considered perfect power-laws throughout.) Possible curvature in integral source counts ${\rm d}N/{\rm d}\Omega(S>S_{*})$ may warrant small corrections to $\tilde{x}$, as illustrated by the difference between the blue and the black curves in Fig. 1. Also the domain over which to integrate $\alpha$, illustrated by the white shaded area in Fig. 2, offers scope for such corrections. Both are expected to be small corrections to an already small effect of $\mathcal{O}(\beta)$, but can nonetheless be estimated directly from data. Related corrections were proposed by Tiwari et al. (2014) and implemented by Siewert et al. (2021). While they did not recognise the adequacy of performing the fit of $\tilde{x}$ only locally—e.g.  Siewert et al. (2021) say their fit “extends for each survey over one decade”—they demonstrated the feasibility of the procedure, albeit with little effect on their conclusions.
Looking ahead, this work showed that the expected kinematic matter dipole amplitude $\mathcal{D}_{\rm kin}$ in general is a function of the flux threshold $S_{*}$, since $\tilde{x}$ and $\tilde{\alpha}$ themselves depend on $S_{*}$, opening up much ground for study (e.g., Murray (2022) used this observation to constrain the lensing dipole contribution). Most importantly however, the present formalism allows to make practical connections between theoretical calculations of the redshift-dependent kinematic matter dipole and the one derived from observed quantities. While measuring the matter dipole’s redshift dependence naturally requires the careful definition of redshift-bins in data, the theoretical computation of its kinematic expectation necessitates equally careful treatment of boundary terms, especially for narrow redshift bins with sharp cut-offs. These terms naturally vanish in samples as those considered here and in DB22 (see also Domènech et al., 2022). Nevertheless, future extension of the present work to accommodate redshift-tomographic measurements of the matter dipole is key to building a full understanding of the matter dipole anomaly. Given that correspondingly suitable galaxy catalogues from current surveys are in reach this could not be more timely.